The Drogon Co. just issued a dividend of $2.51 per share on its common stock. The company is expected to maintain a constant 6 percent growth rate in its dividends indefinitely. If the stock sells for $40 a share, what is the company's cost of equity?

13.28%

12.65%

6.8%

12.28%

12.02%
Answer #1
Solution:
Answer is 2nd option 12.65%
Working Notes:
We will compute cost of equity of the company using constant growth method of Gordon know as Gordon frowth model.
Using Gordon growth model : P0 = D0(1+g) / (Ke - g)
ke = cost of Equity =??
Po=current share price = $40 per share
g=constant growth rate= 6%
D0= Just paid Dividend=$2.51 per share
P0 = D0(1+g)/(Ke -g)
Ke = (D0(1+g)/P0) + g
Ke = (2.51 x (1+6%)/40) + 6%
Ke = 6.6515% + 6%
Ke = 12.6515%
Ke = 12.65%
Hence cost of equity of the company is 12.65%
Please feel free to ask if anything about above solution in comment section of the question.